OpenStudy (anonymous):

Credit Cards. The balance in Will's Mastercard account grew 2%, to $870, in one month. What was his balance at the beginning of the month?

OpenStudy (anonymous):

the balance at the beginning=x interset 2%* x

OpenStudy (anonymous):

x+0.02x=870 x(1+0.02)=870 1.02x=870 x=870/1.02